# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v1

@@ 3. $\bullet$
A common feature of these results is that they provide strong $L^{p}$$L^{q}$ bounds only away from the critical line, while critical-line estimates appear to be unavailable from literature. One main reason is that the strong-type estimate on the critical line generally fails (see, Figure 2). One main reason is that the strong-type estimate on the critical line generally fails (see Figure 2). This obstruction, in turn, makes it difficult to use the techniques that are effective in the cases $t\leq 1$, as $K_{2t}$ becomes more singular in the hypersingular regime (see further discussion after (1.3)).
@@
where $\mathbb{T}$ denotes the unit torus and $Q_{I}$ is the Carleson box associated with an arc $I\subseteq\mathbb{T}$. It is also convenient to introduce its dyadic analogue. Let $\mathcal{D}$ be a dyadic system on $\mathbb{T}$. For $t>1$, the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$ is defined by
@@
    Likewise, if one replaces $\mathbb{D}$} by the upper half-plane, then a careful inspection of the argument in (1.2) yields that the operator is again not well-defined on nonzero constant functions whenever $t>1$.
@@
To address this issue, we develop a new approach based on Bourgain’s interpolation trick (see Lemma 4.2).
@@
A careful examination of the proof of the above estimates for $K_{2t}$ shows that the same bounds also hold for the associated positive operator $K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w)$, which can be regarded as a hypersingular analogue of the Berezin transform.
@@
${\mathcal{S}}_{m}:=\left\{I_{k}:=\left[\frac{k}{2^{m}},\frac{k+1}{2^{m}}\right),0\leq k\leq 2^{m}-1\right\}\cup[0,2).$
@@
We will prove these estimates in Theorem 5.6. For the reader’s convenience, we summarize the resulting bounds below; see Figure 3.
@@ 2. Notations
We say that $\mathcal{T}$ is of weak type $(p,q)$, and write $\mathcal{T}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$, if there exists $C>0$ such that for all $f\in L^{p}({\mathbb{D}})$ and all $\lambda>0$,
@@
Moreover, we say that $\mathcal{T}$ is of restricted weak type $(p,q)$ if $\mathcal{T}:L^{p,1}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$, or equivalently (a well-known fact), if there exists $C>0$ such that for every measurable set $E\subset{\mathbb{D}}$ with $|E|<\infty$ and all $\lambda>0$,
$\big|\{z\in{\mathbb{D}}:\ |\mathcal{T}\mathbf{1}_{E}(z)|>\lambda\}\big|\leq C\,\lambda^{-q}|E|^{q/p},$
@@ 3. $L^{p}$ theory for the hypersingular maximal operator
    Off-critical line regime, that is, when $\left(\frac{1}{p},\frac{1}{q}\right)\in\left\{[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\}$;
@@
    Critical line regime, that is, when $\left(\frac{1}{p},\frac{1}{q}\right)\in\left\{[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\}$.
@@
$\simeq\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}\int_{0}^{1}\frac{dr}{(1-r)^{\frac{2(t-1)}{2(t-1)+\varepsilon}}}$
$\lesssim\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}.$
@@
$\left(\frac{1}{p},\frac{1}{q}\right)\in\left\{[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\},$
@@ 3.2. Critical line estimate
$\sum_{i=1}^{\infty}\sum_{J\subseteq I_{i},\;J\;\textrm{dyadic}}\frac{1}{|Q_{i}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}}$ $\simeq\sum_{i=1}^{\infty}\sum_{k\geq 0}\sum_{\begin{subarray}{c}J\subseteq I_{i},\;J\;\textrm{dyadic}\\ |J|=2^{-k}|I_{i}|\end{subarray}}\frac{1}{2^{k}|Q_{J}^{\textnormal{up}}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}$
@@
$\left(\frac{1}{p},\frac{1}{q}\right)\in\left\{(0,1)^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\},$
then for every $0<r\leq\infty$, the operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ extends a bounded operator
${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p,r}({\mathbb{D}})\to L^{q,r}({\mathbb{D}}).$
In particular, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
@@ 4. Critical line estimates for the hypersingular Bergman projection
Our next goal is to prove estimates at the other endpoint $\big(\frac{1}{3-2t},1\big)$. As discussed in Section 1 (see the comments following (1.3)), the main difficulty is the lack of strong-type bounds along the critical line 1/q1/p=2t21/q-1/p=2t-2. To overcome this difficulty, we combine a “hypersingular” sparse domination with Bourgain’s interpolation trick. (see, e.g., [3, 6]; see also [7] for an abstract extension of this method developed in the setting of fairly general normed vector spaces). The following formulation is the one needed for our purposes.
@@
$\left|K_{2t}f(z)\right|=\left|\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}dA(w)\right|\lesssim{\mathcal{A}}^{t}_{{\mathcal{S}},{\mathcal{D}}}f(z)+{\mathcal{A}}^{t}_{{\mathcal{S}},\widetilde{{\mathcal{D}}}}f(z),\quad z\in{\mathbb{D}},$
@@ 5.1. Graded family
      $p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\qquad\textrm{and}\qquad q=1.$
@@
Theorem 5.6 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see Section 4). Moreover, the proof of Step II in Theorem 5.6 provides an alternative approach to [12, Theorem 3].
@@ 6. Some open problems
1. Upgrading restricted weak type to weak type at the endpoint. It would be interesting to revisit the endpoint estimates for $K_{2t}$ at
$p=\frac{1}{3-2t}\quad\text{and}\quad q=1,$
and for $\mathbb{A}_{{\mathcal{S}}}^{t}$ at
$p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\quad\text{and}\quad q=1.$
We ask whether one can upgrade the restricted weak-type bounds at these points to weak-type estimates.

### v2

@@ 3. $\bullet$
A common feature of these results is that they provide strong $L^{p}$$L^{q}$ bounds only away from the critical line, while estimates on the critical line itself appear to be unavailable in the literature. One main reason is that strong-type bounds on the critical line generally fail (see, Figure 2). This obstruction, in turn, make it difficult to use the techniques that are effective in the case $t\leq 1$, since $K_{2t}$ becomes more singular in the hypersingular regime (see the discussion after (1.3)). Moreover, for general hypersingular operators, weak-type bounds may even fail at one of the endpoints on the critical line (see Remark 1.7), indicating that genuinely new ideas are required to treat critical-line estimates.
@@
where $\mathbb{T}$ denotes the unit circle and $Q_{I}$ is the Carleson box associated with an arc $I\subseteq\mathbb{T}$. It is also convenient to introduce its dyadic analogue. Let $\mathcal{D}$ be a dyadic system on $\mathbb{T}$. For $t>1$, the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$ is defined by
@@
    Likewise, if one replaces $\mathbb{D}$} by the upper half plane $\mathcal{H}:=\{z\in\mathbb{C}:\textnormal{Im}z>0\}$, then a careful inspection of the argument in (1.2) yields that the operator is again not well-defined on nonzero constant functions whenever $t>1$ (see also [12, Theorem 5] for a related statement concerning the hypersingular Bergman projection $K^{\mathcal{H}}_{2t}$ on the upper half plane).
@@
To address this issue, we develop a new approach based on Bourgain’s interpolation trick (see, Lemma 4.3).
@@
  1. (1)
    A careful examination of the proof of the above results for $K_{2t}$ shows that the same bounds also hold for the associated positive operator
    (1.4) $K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w),$
    which can be regarded as a hypersingular analogue of the Berezin transform.
  2. (2)
    One may ask whether, at the endpoint (1/p,1/q)=(32t,1)(1/p,1/q)=(3-2t,1), the restricted weak-type estimate (1.3) can be upgraded to a weak-type bound (recall that $\mathcal{M}_{t}^{\mathcal{D}}$ does satisfy a weak-type estimate at this endpoint). The situation, however, is more subtle. Indeed, one can show that the hypersingular Berezin transform (1.4) is unbounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$ (see, Proposition 4.2). For $K_{2t}$, we conjecture that the weak-type estimate also fails at this endpoint (see Section 6 for further discussion).
@@
${\mathcal{S}}_{m}:=\left\{I_{k}:=\left[\frac{k}{2^{m}},\frac{k+1}{2^{m}}\right),0\leq k\leq 2^{m}-1\right\}\cup\left\{[0,2)\right\}.$
@@
    However, in general the endpoint weak-type estimate for $\mathbb{A}_{{\mathcal{S}}}^{t}$ fails. Indeed, this follows from Proposition 4.2 together with the fact that $K_{2t}^{+}$ can be bounded above pointwisely by a sum of two hypersingular sparse operators (see (4.1)).
We will prove these estimates in Theorem 5.6. For the reader’s convenience, we summarize the resulting bounds below (see Figure 3).
@@ 2. Notations
We say that $\mathcal{T}$ is of weak type $(p,q)$, and write $\mathcal{T}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded, if there exists $C>0$ such that for all $f\in L^{p}({\mathbb{D}})$ and all $\lambda>0$,
@@
Moreover, we say that $\mathcal{T}$ is of restricted weak type $(p,q)$ if $\mathcal{T}:L^{p,1}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded, or equivalently (a well-known fact), if there exists $C>0$ such that for every measurable set $E\subset{\mathbb{D}}$ with $|E|<\infty$ and all $\lambda>0$,
$\big|\{z\in{\mathbb{D}}:\ |\mathcal{T}\mbox{1\hskip-4.25pt{1}}_{E}(z)|>\lambda\}\big|\leq C\,\lambda^{-q}|E|^{q/p},$
@@ 3. $L^{p}$ theory for the hypersingular maximal operator
    Off-critical line regime, that is, when $(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\}$;
@@
    Critical line regime, that is, when $(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\}$.
@@
$\simeq\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}\int_{0}^{1}\frac{dr}{(1-r)^{\frac{2(t-1)}{2(t-1)+\varepsilon}}}\lesssim\left\|f\right\|^{\frac{1}{2t-2+\varepsilon}}_{L^{\infty}}.$
From the above arguments, it is not hard to see that ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ is unbounded from $L^{\infty}({\mathbb{D}})$ to $L^{\frac{1}{2t-2}}({\mathbb{D}})$. ∎
@@
$(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}>2t-2\right\},$
@@ 3.2. Critical line estimate
$\sum_{i=1}^{\infty}\sum_{J\subseteq I_{i},\;J\;\textrm{dyadic}}\frac{1}{|Q_{i}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}$ $\simeq\sum_{i=1}^{\infty}\sum_{k\geq 0}\sum_{\begin{subarray}{c}J\subseteq I_{i},\;J\;\textrm{dyadic}\\ |J|=2^{-k}|I_{i}|\end{subarray}}\frac{1}{2^{k}|Q_{J}^{\textnormal{up}}|^{\frac{1}{2}}}\int_{Q_{J}^{\textnormal{up}}}\frac{dA(z)}{\omega^{\frac{3-2t}{2t-2}}(z)}$
@@
$(p,q)\in\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in(0,1)^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\},$
then for every $0<r\leq\infty$, the operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ extends a bounded operator from $L^{p,r}({\mathbb{D}})$ to $L^{q,r}({\mathbb{D}})$. In particular, ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}({\mathbb{D}})\to L^{q,\infty}({\mathbb{D}})$ is bounded.
@@ 4. Critical line estimates for the hypersingular Bergman projection
Our next goal is to prove estimates at the other endpoint $\big(\frac{1}{3-2t},1\big)$. To begin with, we have the following unboundedness result.
Proposition 4.2.
Let $1<t<3/2$. Then the hypersingular Berezin transform $K_{2t}^{+}$ fail to be bounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$, where we recall that
$K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w).$
Proof.
Take any $\varepsilon>0$ sufficiently small such that $3-2t-\varepsilon>0$, and let $f(z):=\frac{1}{(1-|z|^{2})^{3-2t-\varepsilon}},\;z\in{\mathbb{D}}$. First, by a direct computation, we have
$\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}})}\simeq\frac{1}{\varepsilon^{3-2t}}.$
On the other hand, by [29, Theorem 1.12], one has
$K_{2t}^{+}f(z)$ $=\int_{{\mathbb{D}}}\frac{1}{(1-|w|^{2})^{3-2t-\varepsilon}|1-z\overline{w}|^{2t}}dA(w)$
$=\int_{{\mathbb{D}}}\frac{(1-|w|^{2})^{2t+\varepsilon-3}}{|1-z\overline{w}|^{2t}}dA(w)$
$\simeq\frac{1}{(1-|z|^{2})^{1-\varepsilon}}.$
Therefore, for any $\alpha>0$, the above estimate yields
$\alpha\left|\left\{z\in{\mathbb{D}}:\left|K_{2t}^{+}f(z)\right|\gtrsim\alpha\right\}\right|$ $\simeq\alpha\left|\left\{z\in{\mathbb{D}}:\frac{1}{(1-|z|^{2})^{1-\varepsilon}}>\alpha\right\}\right|$
$\simeq\alpha\cdot\alpha^{-\frac{1}{1-\varepsilon}}=\alpha^{\frac{\varepsilon}{1-\varepsilon}}.$
Thus, if $K_{2t}^{+}:L^{\frac{1}{3-2t}}({\mathbb{D}})\to L^{1,\infty}({\mathbb{D}})$ is bounded, then for any $\alpha>0$, one has
$\alpha^{\frac{\varepsilon}{1-\varepsilon}}\lesssim\frac{1}{\varepsilon^{3-2t}},$
which is impossible by letting $\alpha\to\infty$. ∎
Now we turn to some positive results. As discussed in Section 1 (see the comments following (1.3)), the main difficulty is the lack of strong-type bounds along the critical line 1/q1/p=2t21/q-1/p=2t-2. To overcome this difficulty, we combine a “hypersingular” sparse domination with Bourgain’s interpolation trick. (see, e.g., [3, 6]; see also [7] for an abstract extension of this method developed in the setting of fairly general normed vector spaces). The following formulation is the one needed for our purposes.
@@
(4.1) $\left|K_{2t}f(z)\right|=\left|\int_{{\mathbb{D}}}\frac{f(w)}{(1-z\overline{w})^{2t}}dA(w)\right|\lesssim{\mathcal{A}}^{t}_{{\mathcal{S}},{\mathcal{D}}}f(z)+{\mathcal{A}}^{t}_{{\mathcal{S}},\widetilde{{\mathcal{D}}}}f(z),\quad z\in{\mathbb{D}},$
@@ 5.1. Graded family
      (5.3) $p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\qquad\textrm{and}\qquad q=1.$
@@
Theorem 5.6 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.6 yields an alternative proof of [12, Theorem 3]. Finally, we note that the endpoint weak-type estimate (5.3) fails in general (see, Proposition 4.2).
@@ 6. Some open problems
1. Upgrading restricted weak type to weak type at the endpoint. It would be interesting to revisit the endpoint mapping properties of $K_{2t}$ at
$p=\frac{1}{3-2t}\quad\text{and}\quad q=1.$
We ask whether the restricted weak-type bound at this exponent can be upgraded to a weak-type estimate. We expect that any such refinement, if true, would require genuinely complex-analytic input: one would need to exploit additional cancellation coming from the kernel $(1-z\overline{w})^{-2t}$, rather than working only with the positive kernel $|1-z\overline{w}|^{-2t}$. The latter corresponds to the hypersingular Berezin transform $K_{2t}^{+}$, which fails to be bounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$ (see, Proposition 4.2).